OpenStudy (anonymous):

Please Help!!! The value of Jamie's baseball card collection rose 50% in the last year. It is now worth $240. What was it worth a year ago? $190 $160 $110 $155

OpenStudy (anonymous):

actually B

OpenStudy (anonymous):

Since this is multiple choice, the easiest way to go would be to just find 50% of each choice and add it to said choice until you reach one that equals $240. \[190*.5=95\]\[95+190=285\]and so on. In this case, the answer would be 160 dollars. Half of 160 is 80, and 80 plus 160 = 240, which is the target sum. Therefore, the answer is B

OpenStudy (anonymous):

Could you help me with this one?? Shanna's collection of antique pottery has increased in value over the past two years. Two years ago, the value of her collection was $1,318. If the value increased 6% during the first year and 5% during the next year, what is the current value? $69.85 $1,466.93 $1,464.43 $1,397.08

OpenStudy (anonymous):

Alright. Since you are given the starting value of the collection, all you have to do is find 6 percent of it, add it on, and then find 5% of the new value. \[1318*.06=79.08\] \[1318 + 79.08 = 1397.08\]This is the value of the collection after the first year. Now we just have to do the same thing with this new value and 5%\[1397.08*.05=69.854\]\[1397.08+69.854=1466.934\]Since we are talking about money, we will round to the hundredths place, giving us a final value of 1466.93

OpenStudy (anonymous):

Thankk you could u hellp me with 2 more please

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Maurice has been recording the price of fuel at a local gas station throughout the week. His data is shown below. Day Price (per gallon) Monday $2.68 Tuesday $2.79 Wednesday $2.85 Thursday $2.79 Friday $2.71 To the nearest percent, what is the percent change in the gas price from Wednesday to Friday? 2% decrease 5% decrease 7% decrease 14% decrease

OpenStudy (anonymous):

Okay, with this one, you need to determine the difference in the prices of the starting and ending value. \[2.85-2.71=.14\]We now know that there was a 14 cent change over the course of those three days. Now all we need to figure out how much (in a percentage) 14 cents is out of $2.85. To do this, all we need to do is divide .14 by 2.85.\[\frac{ .14 }{ 2.85 }=.049122807\] To make this easier to read, we can multiply by 100 so this is not a decimal\[.049122807*100=4.9122807\] 4.9 rounds to 5 percent, which gives us answer choice B

OpenStudy (anonymous):

Thank you i have one more sir!! :) Last week, a total of 4,000 fans attended the Baytown College football game. This week, a total of 3,200 fans attended the Baytown College football game. What is the percent decrease in attendance from last week to this week? 84% decrease 89% decrease 20% decrease 80% decrease

OpenStudy (anonymous):

Okay, this question is similar to the first in that we need to determine the difference between the two values and then find the percentage of that difference out of the initial value. \[4000-3200=800\]\[\frac{ 800 }{ 4000 }=.2\]This is clearly a 20% decrease because the initial value is greater than the latter, and the decimal is .2, which is equal to 20/100 (You had better become a fan of me lol)

OpenStudy (anonymous):

Thanky you for all youre help!!!

OpenStudy (anonymous):

You're welcome!